OpenStudy (anonymous):

2x + 3y = 2 x + 6y = 4

OpenStudy (anonymous):

multiply the second one by \(-2\) and then add first step is \[2x+3y=2\\ -2x-12y=-8\] at least that is one way to do it

OpenStudy (anonymous):

2x + 3y = 2 x + 6y = 4 Muliply second equation by -2. 2x + 3y = 2 -2x - 12x =-8 Add the two equations. -9y = -6 y = 2/3 Substitute y into the original equation. 2x + 3(2/3) = 2 2x + 2 = 2 - 2 -2 ------------ 2x = 0 x = 0 So, x = 0 y = 2/3